# PlanQA: A Benchmark for Spatial Reasoning in LLMs using Structured Representations

### v1

@@ 3.3 Evaluation Protocol and Scoring
Each question in PlanQA is paired with a reference answer derived entirely from the symbolic layout, allowing deterministic and automated evaluation of model outputs. Depending on the response type, correctness is assessed using tolerance-based numeric comparison, symbolic string matching, or geometric validation routines. The scoring logic referenced by each question format code is detailed in Table 4, and is applied uniformly across models.
For numerical questions (e.g., distances, areas, angles), predictions are accepted if they fall within a fixed relative error of the reference: $\left|\frac{\hat{y}-y^{*}}{y^{*}}\right|\leq\varepsilon$ where $\varepsilon$ is typically 0.02. This threshold is applied uniformly across models, and reflects spatial resolution limits in architectural layouts.
Binary and categorical responses, such as visibility, object fit, or violation detection, are validated against geometry-aware checks. For example, visibility is determined by ray intersection tests over the layout’s free space, while object fit questions are scored by verifying non-overlap and clearance constraints for a candidate placement. Boolean answers must match the reference exactly.
Structured outputs, including bounding boxes and paths, are checked for both format and semantic correctness. Rather than using exact coordinates, we validate these using geometric plausibility conditions—e.g., sufficient overlap with ground truth, or collision-freeness for planned paths. Thresholds for spatial alignment (e.g., bounding box proximity or path deviation) are set conservatively to account for minor lexical variation without over-penalizing near-correct reasoning.
To differentiate reasoning failures from generation or parsing errors, we annotate each model response with a status flag: Correct, Incorrect, Malformed, or Missing. Malformed responses include answers lacking the required structure “Final Answer:” or acceptable variations thereof, or containing ill-typed data. These annotations allow us to separate breakdowns in spatial inference from extraction failures and language model formatting idiosyncrasies.
@@ 3.4 Model Inference Setup
All models are queried using standard chat-based completion APIs, with prompts constructed as described above. We evaluate a mixture of open-weight and proprietary language models across three resource tiers, distinguished by their token limits and relative inference cost. Models in the “reasoning” tier are allocated up to 8192 tokens per completion, enabling them to process longer layout descriptions and multi-step CoT outputs. Other models operate within a 4096-token context window and are evaluated only on the non-reasoning subset of questions.
No model receives fine-tuning or prompt adaptation specific to PlanQA. All prompts are zero-shot, with the system message and output formatting constraints held fixed. For each model, we use the default inference configuration provided by its vendor, including temperature settings. In practice, these are near-deterministic across runs and sufficiently stable for head-to-head comparison. In cases where a model produces variable-length completions or truncates early, we retain the first parseable instance of the “Final Answer:” line, if present.
Each model is evaluated over a batch of 600 layouts per room type, with one question per layout. Requests are grouped into new-line-delimited JSON (JSONL) payloads for efficient submission via batched APIs. Layout serialization, prompt insertion, and output parsing are fully automated, with no manual curation or intervention.
We include both high-capacity generalist models and compact instruction-tuned variants to measure the degree to which spatial reasoning emerges with scale and supervision alone. All models are evaluated on identical input distributions and scoring criteria, enabling cross-system comparisons that are both cost-aware and architecture-agnostic.
@@ 4 Results
We evaluated the performance of the model on a dataset of 1,800 layouts, consisting of 600 kitchens, 600 living rooms and 600 bedrooms. Each layout is paired with one question sampled from a pool of 10 parameterized templates, filtered by room applicability. This yields 18,000 layout–question pairs. The full taxonomy is shown in Table 2.
Eleven language models are evaluated, covering a wide range of parameter scales, architectures, and training regimes. The reasoning-oriented models include Qwen3-32B Yang et al. (2025) and DeepSeek-R1 DeepSeek-AI (2025). The large general-purpose models include DeepSeek-V3 DeepSeek-AI (2024), LLaMA 3.3-70B-Instruct, Gemma 2-27B-it Riviere et al. (2024), phi-4 Abdin et al. (2024b), and GPT-4.1 OpenAI et al. (2024). The smaller models include LLaMA 3.1-8B-Instruct Grattafiori et al. (2024), Gemma 2-9B-it Riviere et al. (2024), Phi-3.5-mini Abdin et al. (2024a), and GPT-4o-mini. All models are evaluated in identical zero-shot conditions using standardized prompts and serialized layout inputs, as described in Section 3. The complete results disaggregated by question, room type, and model are included in the appendix.
@@ 4.1 Quantitative Results
We begin by aggregating accuracy across cognitive targets as defined in Table 4. Figure 2 shows a heatmap of mean accuracy by model and target category. Metric reasoning (M) and visual logic (V) show the highest cross-model scores, while Topology (T), rule-based validation (R), and geometric planning (G) are more difficult across the board. DeepSeek-R1 (reasoning), GPT-4.1 (large), and GPT-4o-mini (compact, tuned) consistently lead the performance in the harder categories, although even large models struggle in certain occlusion and clearance tasks. Surprisingly, GPT-4.1’s performance on vector-direction (V) questions nearly matches DeepSeek-R1 and GPT-4o-mini—despite those models’ explicit chain-of-thought optimization—highlighting an unexpected parity in occlusion reasoning.
Figure 2: Accuracy by model and cognitive target. Each cell shows average accuracy across questions associated with a given skill cluster. Cognitive targets: M = Metric computation, T = Topology, G = Geometric planning, V = Visual logic / occlusion, R = Rule-based validation.
Performance improves with model scale, but not uniformly across tasks. As shown in Figure 3, accuracy improves steeply up to  30B parameters, suggesting that scale alone cannot fully close the gap on complex spatial tasks. Notably, gains largely plateau beyond 30 B parameters, indicating diminishing returns for raw scale on planning and constraint problems.
Figure 3: Accuracy vs. model size. Models are arranged from the smallest to largest in terms of trainable parameters. While larger models perform better overall, instruction tuning and task specialization also contribute significantly.
We also compare accuracy across room types (Figure 4). Kitchens exhibit the highest overall accuracy, likely because their regular counter-and-appliance grids reduce occlusion and simplify clearance reasoning. In contrast, living rooms and bedrooms generally yield lower performance across models, due to denser layouts, complex furniture adjacencies, and higher occlusion likelihood, with frequent object groupings such as televisions on stands, sofas and coffee tables on rugs, or lamps on nightstands. These room-type gaps suggest that layouts with strong grid structure provide easier inductive priors for LLMs than organic, cluttered arrangements.
Figure 4: Accuracy by question and room type. Kitchens show the highest average accuracy, likely due to their regular, grid-like counter layouts and lower occlusion. The inability of most models to estimate free space in living rooms and bedrooms may be caused by more clutter, or the presence of objects like area rugs which the model was instructed to ignores for occlusion purposes.
@@ 4.2 Ablation
To evaluate the robustness of layout interpretation under alternate encodings, we perform a format ablation that replaces the standard JSON layout with a semantically equivalent XML version. This substitution preserves all geometric and object-level content while modifying only the syntax and structural serialization. To reduce evaluation cost, we focused on the most variance-sensitive question categories in the kitchen domain: Repositioning, Vector Direction, and Free Space. These were selected based on standard deviation in accuracy across the full model set. We apply each ablation to three illustrative models spanning architecture classes and capacity tiers: GPT-4.1 (closed large), GPT-4o-mini (compact tuned), and DeepSeek-R1 (reasoning-optimized). As shown in Table 5, accuracy remains stable across formats for most question types, suggesting that these models encode layout semantics in a manner invariant to low-level representation details.
@@
Model Repositioning View Angle Obstruction
GPT-4.1 77.0 → 76.5 92.2 → 93.7 75.5 → 89.6
GPT-4o-mini 33.8 → 29.6 42.4 → 42.2 12.2 → 15.8
DeepSeek-R1 83.6 → 93.7 74.7 → 74.3 95.7 → 96.1
@@ 5 Conclusion
We have presented PlanQA, a benchmark that evaluates spatial reasoning over symbolic 2D layouts by leveraging object-centric encodings (JSON, XML) similar to those used in architecture and robotics. Through a systematic evaluation of 11 language models across five skill clusters: metric computation, topology, geometric planning, visual logic, and rule-based validation, we make critical findings with implications for the use of LLMs in geometric design tasks.
Metric and basic visibility queries (distances, areas, line-of-sight) are handled reliably (>95% accuracy), whereas geometric planning, topological relations, and constraint verification consistently fall below 50% accuracy. Performance gains level off beyond approximately 30 billion parameters, and models optimized for chain-of-thought show only modest improvements on the hardest tasks. Additionally, grid-like layouts (kitchens) yield higher accuracy than more varied configurations (living rooms, bedrooms), indicating that layout regularity influences model performance.
These results reveal that current LLMs do not form sufficiently robust internal representations for complex spatial inference. Addressing this limitation could take two complementary directions. In the near term, integrating external geometric solvers or symbolic planning modules can compensate for the models’ weaknesses in collision avoidance and clearance reasoning. In the longer term, training regimes that incorporate explicit spatial constraints—through specialized objectives, data augmentation with irregular layouts, or constraint-violation examples—may be necessary to improve performance on design-oriented tasks.
@@ Acknowledgments
The work is supported by funding from KAUST, Center of Excellence for Generative AI, under award number 5940, and the NTGC-AI program
Appendix A Technical Appendices and Supplementary Material for PlanQA
This technical appendix includes an overview of the dataset generation process, followed by full evaluation results on our generated QA pairs, along with aggregated metrics and ablation studies. It also presents an extended taxonomy of spatial reasoning tasks, explanations of selected failure cases, an expanded analysis of vision-language models (VLMs), and examples of prompts used for both generation and questioning.
The full dataset and all code used for generation, evaluation, and visualization are provided in the supplementary materials.
@@ Appendix B Details on Synthetic Data Generation
Room layouts were generated using Gemini 2.5 Pro, specifically fine-tuned for spatial reasoning and bounding-box tasks Team et al. (2025). Below, we describe the detailed procedure and constraints. All generation scripts, prompts, constraint-checking code, and random seeds are included in the code for full reproducibility.
@@ B.1 Room Shape Generation
We generated 600 layouts each for kitchens, bedrooms, and living rooms. These layouts include a variety of geometries and sizes, with clearly defined room structures containing windows, doors, and corner cutouts where applicable. All layouts follow specifications tailored to each room type.
Each layout falls into one of three shape categories: rectangular, L-shaped, or open. Size categories were defined individually for each room type and were incorporated into the generation prompts based on common assumptions about typical room dimensions.
@@
Room Type Shape Distribution (Rect / L-shaped / Open) Size Categories (in m2) (Small / Medium / Large)
Kitchen 40% / 40% / 20% $\leq$7 / 7–18 / $>$18
Bedroom 50% / 30% / 20% 8–12 / 12–18 / $>$18
@@
Room geometries were procedurally varied using prompt-based guidance to produce rectangular, L-shaped, and open-plan configurations. The following structural properties were described in the prompts but not explicitly enforced during layout generation:
  • L-shaped rooms were described as rectangular spaces with a square cutout from one corner. To maintain usable proportions, each leg of the L shape was suggested to be at least 1.5,m wide and deep.
  • Open-plan rooms, apart from living rooms, were prompted without doors and with one full wall removed. This was intended to vary room shapes and simplify layout generation.
  • Doors were described with widths between 0.8,m and 1.0,m, selected randomly. Windows were chosen from a fixed set of widths: 0.6,m, 0.75,m, 0.9,m, 1.2,m, and 1.5,m.
  • Prompts included guidance to place all elements such as doors, windows, and cutouts entirely within room boundaries.
@@
  • Windows on the same wall were the same size to support visual balance. Small, isolated windows were not used.
@@
  • Windows were not placed opposite each other or on the same wall as a door, as such arrangements are less common and were not emphasized in the prompt design.
@@ B.3 Layout Selection Criteria
Approximately one-third of the initially generated room layouts were filtered-out using a set of geometric and functional constraints. These filters were designed to ensure realistic object placement, functional usability, and architectural plausibility. While these rules are based on general design principles, they are not based on any specific design standard. Instead, they are the result of iterative development, focusing specifically on addressing cases where our prompts lead to unlikely or implausible layouts. After filtering, we retained 600 valid layouts for each room type.
@@
  • Non-blocking door clearance: Doors have physical thickness and are defined by bounding boxes $(x_{1}^{d},y_{1}^{d},x_{2}^{d},y_{2}^{d})$. A clearance zone of $door\_length$ meters is required in front of the door to ensure swing space and accessibility. The position of this zone depends on which wall the door is attached to.
@@
B.4 Room Statistics
Table 7 summarizes key statistics for the generated layouts across room types, including average object count, overlap frequency, and spatial density. Kitchens typically contain fewer objects but have higher density due to their compact size. In contrast, bedrooms and living rooms host more objects with greater overlap, reflecting more open layouts and flexible furnishing.
This structural simplicity, with fewer objects, minimal overlap, and higher density, makes kitchens the simplest environment for spatial reasoning tasks. Due to the limited amount of free space, some complex questions, such as those involving paths, have fewer possible answers. As a result, models tend to perform better on kitchen-related questions.

### v2

@@ 3.5 Evaluation Protocol and Scoring
Each question in FloorplanQA is paired with a reference answer computed directly from the symbolic layout, enabling fully automated and deterministic evaluation of model outputs. Depending on the response type, correctness is assessed using numeric comparison with fixed tolerances, string matching, or geometric validation checks.
For numerical questions (e.g., distances, areas, angles), predictions are accepted if they fall within a relative error threshold. Sequence outputs are evaluated for both format and semantic correctness. Rather than requiring exact coordinate matches, we validate paths using geometric plausibility conditions of collision avoidance and sufficient proximity to a ground-truth trajectory. Deviation thresholds are set conservatively to tolerate minor geometric variations without crediting qualitatively wrong routes.
To differentiate reasoning failures from extraction or formatting issues, we apply a regex-based parsing pipeline, covering a fixed set of expected answer patterns (e.g., ‘*Final answer*’ tokens in lower case or with surrounding symbols). If no answer is produced, or if the extracted content does not match a valid format, we count the response as an error (which is also considered incorrect). We provide a detailed breakdown on accuracy in Appendix E. In our evaluations, the proportion of invalidly formatted answers is below 1%.
We also explicitly track cases where no answer is returned due to truncation (API: stop_reason = Token Limit), a failure mode that disproportionately affects reasoning-heavy models. To account for this, we report in Appendix I both the percentage of responses truncated by token limits and an adjusted accuracy computed only over valid (non-truncated) answers. These adjusted accuracies can be interpreted as an approximate upper bound on overall dataset performance, since with unlimited token budgets models could in principle reach that quality.
@@ 3.6 Model Inference Setup
All models are queried using standard chat-based completion APIs, with prompts constructed as described above. We evaluate both large and mid-size models, including reasoning and standard variants. Large reasoning-oriented and standard models are allocated up to 12,288 tokens per completion, enabling them to process long layout descriptions and produce multi-step outputs. Mid-size models, including the GPT family variants optimized for speed and Qwen3-30B, are limited to 8,192 tokens. These budgets reflect our observations that larger models generate longer intermediate justifications and thus consume more tokens.
GPT-5 is evaluated under a distinct configuration: its reasoning intensity can’t be disabled, so we run it with reasoning and verbosity set to “low,” with a maximum output length of 4,096 tokens, while GPT-5-mini is run with reasoning and verbosity set to “medium,” with a maximum output length of 8,192 tokens.
No model receives fine-tuning or prompt adaptation specific to FloorplanQA. All prompts are zero-shot, with the system message and output formatting constraints held fixed. For each model, we use the default inference configuration provided by its vendor, with temperature set to 0. The only exception is GPT-5, for which the temperature cannot be modified and defaults to 1.
Each model is evaluated over 1800 generated layouts and 200 layouts from HSSD, with one question from each type posed per layout. Evaluation is fully automated, from layout serialization and prompt insertion to parsing, with no manual curation.
All models are evaluated on identical input distributions and scoring criteria, enabling cross-system comparisons that are architecture-agnostic and directly comparable.
@@ 4 Results
We evaluated the performance of the model on a dataset of 2,000 layouts, consisting of 600 kitchens, 600 living rooms, 600 bedrooms, and 200 additional layouts from HSSD. Each layout is paired with one question sampled from a pool of 8 parameterized templates, filtered by room applicability. This yields 16,000 layout–question pairs. The full taxonomy is shown in Table 2.
We evaluate fifteen models, spanning a wide range of parameter scales, architectures, and training regimes. The reasoning-oriented models include GPT-5 (OpenAI, 2025a), GPT-OSS-120B (OpenAI, 2025b), DeepSeek-R1-0528 (DeepSeek-AI, 2025), GPT-5-mini, Gemini Flash 2.5 (Google DeepMind, 2025), GPT-OSS-20B, and Qwen3-30B-A3B-Thinking-2507 (Yang et al., 2025). The general-purpose (standard) models include Claude Sonnet 4 (Anthropic, 2025), GPT-4.1 (OpenAI et al., 2024), Moonshot Kimi-K2-Instruct (Kimi Team, 2025), Qwen3-Coder-480B-A35B-Instruct, Qwen3-235B-A22B-Instruct-2507, GPT-4.1-mini, Qwen3-30B-A3B-Instruct-2507, and Devstral-Small-2505 (MistralAI, 2025). All models are evaluated in identical zero-shot conditions using standardized prompts and serialized layout inputs, as described in Section 3. Complete results disaggregated by question type, room type, and model are provided in Appendix E.
@@ 4.1 Quantitative Results
We begin by aggregating accuracy across models and question types for both reasoning and general model families. Figure 2 summarizes accuracy: the top row shows general models, and the bottom row shows reasoning models. In each figure, the left panel summarizes accuracy by model, and the right panel summarizes accuracy by question across room types.
Kitchens lead across models because overlaps are rare, so most queries are straightforward. Scores on HSSD tend to lag behind the other room types; irregular, non-axis-aligned geometry and denser overlap make these layouts more demanding, see Appendix D for detailed statistics. Bedrooms and living rooms are mid-tier and nearly equal, lying between kitchens and HSSD.
Comparing model families, general models struggle when many overlaps must be merged; they often treat object areas independently (no union), which hurts Free Space and Max Box and carries over into path planning. Reasoning models handle those cases with unions and rotations better, so they show gains on Free Space and Max Box. A practical limitation remains: Gemini Flash 2.5 and DeepSeek-R1 hit token limits on larger layouts, which drops scores, especially in HSSD.
Task difficulty follows a consistent pattern in terms of accuracy. Metric-category questions, such as Pair Distance and View Angle, achieve the highest accuracies. Repositioning, Visibility, and Placement yield mid-range accuracies. Max Box and Free Space benefit most from reasoning-oriented models, yet their accuracies remain low and are comparable in difficulty to Shortest Path. Accuracy decreases as object count and overlap density increase, most notably on HSSD layouts. Detailed analysis on each question category is provided in Appendix G
Figure 2: Accuracy of general (top) and reasoning (bottom) models. Left: by model. Right: by question. Each column corresponds to a specific room type represented in our dataset: Kitchens, Living Rooms, Bedrooms (synthetic subsets), plus HSSD layouts.
@@ 4.2 Ablation
To evaluate the robustness of layout interpretation under alternate encodings, we perform a format ablation that replaces the standard JSON layout with a semantically equivalent XML version. This substitution preserves all geometric and object-level content while modifying only the syntax and structural serialization. The prompt level and semantic ablations are described in Appendix, Section F. To avoid recomputing the full suite, we focus on the most variance-sensitive question types for HSSD layouts, as shown in Appendix F (Figs. 5 and 6): View Angle, Visibility, and one task from a different category—Repositioning (Dynamic). We apply each input-format ablation to three representative models: two reasoning models (GPT-OSS-120B, GPT-OSS-20B) and one large general model (Qwen3-235B-A22B) for the same subset of HSSD layouts. As shown in Table 3, accuracy is largely stable in both formats for most categories; the GPT-OSS models change minimally, while Qwen shows modest fluctuations. This suggests that these models encode layout semantics in a manner that is relatively invariant to low-level representation details.
@@
Model Repositioning View Angle Visibility
GPT-OSS-120B 60.5 → 59.0 74.0 → 74.0 70.0 → 72.0
GPT-OSS-20B 40.0 → 39.0 37.5 → 38.5 45.5 → 43.5
Qwen3-235B-A22B 39.0 → 42.0 50.5 → 54.0 70.5 → 65.5
@@ 5 Conclusion
We introduced FloorplanQA, a benchmark for spatial reasoning over symbolic 2D layouts aligned with architecture and robotics practice. We evaluated 15 language models (8 general, 7 reasoning) on 2,000 layouts (1,800 synthetic; 200 semi–real HSSD) across tasks ranging from metric queries to visibility, placement, and shortest path with clearance.
Empirically, metric and simple visibility queries are reliable; kitchens score highest because overlaps are rare. HSSD layouts are more demanding—irregular, non–axis-aligned shapes and dense overlaps expose weaknesses such as centroid miscalculation and missed unions. Reasoning models improve notably on Free Space and Max Box by handling overlaps and rotations more consistently, while general models often subtract object areas independently and fail under heavy overlap. Shortest Path is sufficiently challenging because it requires multiple correct steps (clearance buffering, collision checks, and path search), where errors compound.
These results indicate that current LLMs lack sufficiently robust internal geometric representations for complex spatial inference. Two complementary directions follow. Near term: hybridize with external geometric solvers or symbolic planning modules—set operations (unions/differences), centroid via shoelace, clearance buffering, oriented rectangle search, and A* path planning—to compensate for the models’ weaknesses in collision avoidance and clearance reasoning. Longer term: train with explicit spatial constraints and harder distributions (irregular, overlap–heavy layouts), and include constraint–violation exemplars and geometry-aware objectives so models learn to maintain coherence under rotation, clearance, and union operations in design-oriented tasks.
@@ Acknowledgments
The work is supported by funding from King Abdullah University of Science and Technology (KAUST)—Center of Excellence for Generative AI, under award number 5940, and the NTGC-AI program
Appendix A Technical Appendix and Supplementary Material for FloorplanQA
This appendix details the dataset generation pipeline, dataset statistics, and the full evaluation of FloorplanQA across reasoning and general models. It also includes ablation studies, analysis of failure cases, and the exact prompts used for data generation and question asking.
All artifacts (code, generation prompts, evaluation scripts, and visualization utilities) are included with the submission as supplementary material.
@@ Appendix B Details on Synthetic Data Generation
Room layouts were generated using Gemini 2.5 Pro, specifically fine-tuned for spatial reasoning and bounding-box tasks. In the following sections, we describe the detailed procedure and its constraints. All generation scripts, prompts, constraint-checking code, and random seeds are included in the code for full reproducibility.
@@ B.1 Room Shape Generation
We generated 600 layouts for each of the kitchens, bedrooms, and living rooms. These layouts feature a range of geometries and sizes, with clearly defined room structures that incorporate windows, doors, and corner cutouts where applicable. All layouts follow specifications tailored to each room type.
Each layout falls into one of three shape categories: rectangular, L-shaped, or open. Size categories were defined individually for each room type and were incorporated into the generation prompts based on standard assumptions about typical room dimensions.
@@
Room Type
Shape Distribution
(Rect / L-shaped / Open)
Size Categories (in m2)
(Small / Medium / Large)
Kitchen 40% / 40% / 20% $\leq$7 / 7–18 / >>18
Bedroom 50% / 30% / 20% 8–12 / 12–18 / >>18
@@
The geometry of the room was procedurally varied using prompt-based guidance to produce rectangular, L-shaped, and open-plan configurations. The following structural properties were described in the prompts but not explicitly enforced during layout generation:
  • L-shaped rooms were described as rectangular spaces with a square cutout in one corner. To maintain usable proportions, each leg of the L shape was suggested to be at least 1.5 m wide and deep.
  • Open-plan rooms, apart from living rooms, were prompted without doors and with one whole wall removed. This was intended to vary the room’s shape and simplify the layout generation process.
  • Doors were described with widths between 0.8 m and 1.0 m, randomly selected. The windows were chosen from a fixed set of widths: 0.6 m, 0.75 m, 0.9 m, 1.2 m, and 1.5 m.
  • The prompts included instructions on placing all elements, such as doors, windows, and cutouts, entirely within the room boundaries.
@@
  • The windows on the same wall were the same size to support visual balance. Small, isolated windows were not used.
@@
  • The windows were not placed opposite each other or on the same wall as a door, as such arrangements are less common and were not emphasized in the prompt design.
@@ B.3 Layout Selection Criteria
Approximately one-third of the initially generated room layouts were filtered out using a set of geometric and functional constraints. These filters were designed to ensure realistic object placement, functional usability, and architectural plausibility. While these rules are based on general design principles, they are not based on any specific design standard. Instead, they are the result of iterative development, focusing specifically on addressing cases where our prompts lead to unlikely or implausible layouts. After filtering, we retained 600 valid layouts for each room type.
@@
  • Non-blocking door clearance: Doors have physical thickness and are defined by bounding boxes $(x_{1}^{d},y_{1}^{d},x_{2}^{d},y_{2}^{d})$. A clearance zone of door_lengthdoor\_length meters is required in front of the door to ensure swing space and accessibility. The position of this zone depends on which wall the door is attached to.
@@
Appendix C Details on HSSD layouts selection
We curate a subset of HSSD layouts to ensure clean geometry and unambiguous supervision for spatial reasoning tasks. Starting from the raw scenes, we generate a 2D floor-plan projection and apply filtering and normalization steps (e.g., geometry cleanup and category unification).
Object filtering.
To reduce visual clutter and retain only objects essential for spatial reasoning, we exclude purely decorative or small accessory categories. The banned labels are: accessory, air conditioner, artwork, blanket, boots, bottle, bowl, box, book, brush, candle, cushion, decor, dog, fan, flower, frame, guitar, herb, hook, jar, light, lightbulb, orchids, pendant, pendulum, pet, some plants, poster, pot, shoes, socket, switch, vase, wreath. We also drop subcomponents that fragment footprints without changing free space (e.g., chair/table legs; bed/armchair frames), remove partial cabinet-door leaves, and consolidate multiple near-duplicate small items by keeping a single representative instance. The resulting scenes preserve the functional layout while simplifying geometry. In Figure 3, the left panel shows the layout immediately after 2D projection; the right panel shows the simplified layout after filtering and normalization.
Projection and cleanup.
During 2D projection, we correct mislabeled windows and doors, snap nearly collinear edges, and resolve self-intersections. To smooth rectilinear artifacts and irregular edges in box-like footprints, we apply alpha-shape–based convex hulls, then recompute centroids and areas on the cleaned geometry.
Figure 3: HSSD layout selection: comparison of the original (left) and simplified (right) 2D layouts. The simplified version removes redundant or overlapping objects and cleans geometry to produce well-structured input for spatial reasoning tasks.
Appendix D Room Statistics
Table 5 reports summary statistics for the layouts. Kitchens are typically the smallest spaces, with relatively few objects and overlaps, but a high density due to their compact geometry. Living rooms are the largest, with slightly more objects overall but lower density, reflecting their open layout. Bedrooms fall in between, with similar object counts to kitchens but more frequent overlaps.
The HSSD layouts are comparable in scale to bedrooms and living rooms in terms of area and object count, but differ in structure: objects are represented with detailed, non-axis-aligned polygons, resulting in a significantly higher vertex count. They also exhibit more overlaps than the generated layouts, reflecting their closer alignment with human-authored floorplans. In other respects, however, the distributions remain broadly consistent.